OpenStudy (anonymous):

? over 8 = 16 over 64

OpenStudy (anonymous):

2/8 = 16/64 = 8*2/8*8

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[2 \over 8 = 16\over64\]

OpenStudy (anonymous):

what percent of the graph is shaded ?

OpenStudy (anonymous):

27% of the graph is shaded and the ratio is 27:100

OpenStudy (anonymous):

x=2